OpenStudy (anonymous):

–3x + 2y = 8 change to slope intercept form show all steps.

OpenStudy (anonymous):

Slope intercept form is y=mx+b Step 1: Isolate the term with the y variable

OpenStudy (anonymous):

2y=11

OpenStudy (anonymous):

You can't combine 3x and 8 because 3x has a variable and 8 doesn't

OpenStudy (anonymous):

ok so?

OpenStudy (anonymous):

It will be 2y=3x+8, but you still have to remove the number attached to y by dividing both sides by that number

OpenStudy (anonymous):

how do you do that

OpenStudy (anonymous):

@hal_stirrup

OpenStudy (anonymous):

can you help ?

OpenStudy (anonymous):

Don't you know how to manipulate equations ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

then you should go back to basic equations and practice that so you can get this. dont build a house of card . solid maths foundation is key.

OpenStudy (anonymous):

You need to divide both sides by 2 to remove the 2 that is with the y variable If it was a fraction, you would multiply by the denominator to cancel it out

OpenStudy (anonymous):

For the other side, simply divide both terms by 2 to get your answer.

OpenStudy (anonymous):

2y=3x+8 3x 4y ??

OpenStudy (anonymous):

ok . lets say y/2=x solve for x? what operator do you need to use?

OpenStudy (anonymous):

idkk :(

OpenStudy (anonymous):

think about it.

OpenStudy (anonymous):

divide?

OpenStudy (anonymous):

divide what?

OpenStudy (anonymous):

y and 2?

OpenStudy (anonymous):

nop

OpenStudy (anonymous):

dammit

OpenStudy (anonymous):

ok if i say to you make y/2 to y what do you have to do. note that y/2 is really \[y\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

darn idk

OpenStudy (anonymous):

Sh!+

OpenStudy (anonymous):

its easy . trust me. ;)

OpenStudy (anonymous):

Can you just please give me the answer? this is the last question.

OpenStudy (anonymous):

no sorry NOT allowed . website rules.

OpenStudy (anonymous):

rules are mint to be broken

OpenStudy (anonymous):

JK

OpenStudy (anonymous):

no not here and not for me.

OpenStudy (anonymous):

Alright sounds like people are harsh over here. The slope formula is mx+b=y So thus you need to get the equation above like it.. -3x+2y=8 First, I would add 3x to both sides. So it's like 3x+(-3x) +2y = 8 -3x, the 3x's on the left will cancel. So you will get 2y = 8-3x. So now it's almost in the mx + b =y form you have have the extra 2 in the 2y So now you need to divide both equations by 2 to eliminate the 2y. Thus it would be like... 2y/2 = 8-3x/2 or you can write it like y = 8/2-3x/2... finally the 8/2 can be reduced to 4 SO finally the answer would be... y = 4 - 3/2x

OpenStudy (anonymous):

actually time out let me fix that.. it won't be -3/2x it should be +3/2x it will actually be y = 4 + 3/2x

OpenStudy (anonymous):

Thank you very much

OpenStudy (anonymous):

You're welcome!